OpenStudy (anonymous):

Factorize x^4-11x^2-1

OpenStudy (anonymous):

sory its x^4-11x^2+1

OpenStudy (campbell_st):

sorry I don't think it factorises.... may be a typo

OpenStudy (anonymous):

It is a quadratic equation where x^2 = y. Use the quadratic formula to figure the roots.

OpenStudy (campbell_st):

but then you have u^2 - 11u + 1 which won't factorise

OpenStudy (anonymous):

don't rely only on "factorizing" by ease. Quadratic formula does generate factors you know.

OpenStudy (akshay_budhkar):

it will factorize, \[b^2-4ac >0\]

OpenStudy (campbell_st):

lol..this is strange as the solution is \[(x + (11 +\sqrt{117})/2) (x + (11 - \sqrt{117})/2)\] lol...

OpenStudy (campbell_st):

glad its not my class

OpenStudy (anonymous):

:) Not all factors are "clean as a whistle integers". In real life, they are even worse.